Tinky WInky is 8 years older than Po. Four times Po's age less seven times Tinky Winky's age is 2. How old are they? I have tried the formulas P=T-8 & 4P-7T=2, which doesn't work.
Answer by josgarithmetic(39618)

w for Tinky Winky's age
p for Po's age
-
First sentence translated is w=8+p.
Second sentence translated is 4p-7w=2, but this "less" part seems unclear to me. It might really mean, 7w-4p=2. I need to see a "than" in use before I can understand exactly.
-
YOU said that the one of mine, 4p-7w=2 does not work. MAYBE my other one does?

Substituting for w into what might be the proper needed second equation:
7(8+p)-4p=2
56+7p-4p=2
3p+56=2
Which neither works!

Trying the other "second" equation instead:
4p-7(8+p)=2
4p-56-7p=2
-3p-56=2
3p+56=-2
3p=-58
Which does not work, just as you had found.

I find no other interpretation or translation for that second sentence.
-----WAIT! Maybe there IS another!
-----
-----
4(Po -7w) is 2. Still unclear because missing "than". Maybe one of these then:
4(p-7w)=2 OR 4(7w-p)=2.
Let me try one:
4p-28w=2, substitute,
4(w-8)-28w=2
4w-32-28w=2
-24w-32=2
24w+32=-2
We see this will NOT work.

Last possibility:
28w-4p=2, substitute,
28w-4(w-8)=2
28w-4w+32=2
24w=2-32
Which obviously will only give a w<0, meaning the equation did not work either.